2. a) Use the Laplace transform method to solve: y + y + 1.25 y = t - UT y(0) = 0, y(0) = 0. b) Draw a careful graph of th
Answer #1

(a) given that, Taplace transform method. + (0)20 V tx+1.254=t-ut (+) (+- 7 / 2), y(o)= 0 Taking on both side laplace transfdefine the foreing function for that part we have to est f(t) = -4/2 (t/h) its $4) o, 024<th t +T/ A t, ost</ sto ) (15). ()